solve each system of equations by substitution help i cant figure this out i have the formulas but dont seem to work out right
5x+10y=20
2x+6y=10
6y=-2x+10
y=-1/3x +5/3
5x+10(1/3x+5/3)=20
5x+

Answer by mananth(16949)

5 x + 10 y = 20 .............1
2 x + 6 y = 10 .............2
Eliminate y
multiply (1)by -3
Multiply (2) by 5
-15 x + -30 y = -60
10 x + 30 y = 50

Add the two equations

-5x=-10

/-5

x=2

plug value of x in (1)

5x+10y=20
10+10y= 20
10y=20-10
10y=10
y=1

CHECK

5x+10y= 20
10+10 =20
20 = 20
The rest of them are to be solved in the same way